OpenStudy (anonymous):

Find the total area of the regions enclosed by the curve y=x^4-x^3, the x axis and the line x=2. the answer should be 2.5

OpenStudy (amistre64):

integrate from y from 0 to 2, but you might want to check for any x intercepts along the interval

OpenStudy (amistre64):

y factors as y = x^3 (x-1) y = 0 when x=0 and x=1, so you will simply have two integrations to deal with

OpenStudy (amistre64):

1/20 + 49/20 = 50/20, 5/2 since it asks for the area, area is always positive. otherwise we would simply integrate from 0 to 2 and get about 2.4 which is a measure of displacement.